# Integrated modeling of boron powder injection for real-time plasma-facing component conditioning

### v1

@@ 6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [30]) are shown in figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s. The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-1s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-1s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10
@@
Adjusting the boron injection rate modifies the surface composition of divertor targets, as shown in Figure 6. With a boron influx rate of $10^{20}$ B/s, the B surface concentrations on the inner and outer targets reach approximately 35%, with boron area densities ranging from $0.3-1.4\times 10^{16}$ cm-2. Increasing the influx rate tenfold to $10^{21}$ B/s significantly enriches the boron on the target surfaces, potentially forming nearly pure boron layers, with area densities between $1.8$ and $3.6\times 10^{16}$ cm-2. This saturation phenomenon contrasts with lower influx rates, where the surface composition sustains a mixed B-C coating, approximating a B concentration closer to experimental findings (with B, C concentrations each being 40%). This effect can be attributed to the differential sputtering rates of B and C, which minimize the overall sputtering yields. This effect and associated boron-carbon compositions have been observed in previous DIII-D IPD experiments and traditional boronization procedures in tokamaks like TEXTOR and ASDEX-Upgrade[2, 31, 32]. For IPD experiments, however, achieving a balanced B:C concentration typically required B powder rates of 10-20 mg/s or $10^{21}$ B/s. Discussions in the literature, such as in [2], highlight a marginal dependence of surface composition on boron influx rates, underscoring the challenges of attaining pure boron layers in environments with significant carbon presence. The discrepancies between modeling and experiment may be partially due to the incomplete coverage of the in-vessel wall in the simulation domain. Carbon wall elements beyond the divertor might contribute to the total C influxes due to additional plasma-wall interactions (e.g., main wall CX-erosion) not considered in the present study.
@@ 7 Summary and conclusions
A new comprehensive modeling approach for real-time wall conditioning and in-situ boron powder coating is presented, utilizing three advanced models: EMC3-EIRENE, DIS, and WallDYN3D. These models collectively address 3D plasma transport, powder particle migration and ablation, and mixed-material surface evolution. Ablation modeling with EMC3-EIRENE/DIS shows that boron particles of up to 150 $\mu$m evaporate at the separatrix, where they reach sublimation temperature. The transport analysis shows that initial B fluxes are higher on the inboard divertor target due to the directional frictional forces from the main ion flow in the sublimation region. This effect closely aligns with findings from EMC3-EIRENE simulations using a single atomic B source in the same plasma scenario previously [17], and is consistent with beryllium transport modeling for JET and ITER using ERO2.0 [33]. However, first-flight deposition based on EMC3-EIRENE/DIS alone neglects impurity recycling, resulting in significant asymmetries in the divertor B fluxes concentrated in a narrow, far SOL region. The powder particle mass was not found to affect the distribution of B fluxes between the inboard and outboard divertor targets for particle sizes 5-250 $\mu$m in the present plasma scenario.
@@
This integrated modeling approach provides a new framework for studying real-time wall conditioning and in-situ coating in advanced tokamak scenarios. The results suggest that poloidal and toroidal coverage of plasma-facing components with boron can be achieved through single-point B powder injection within the limitations of the model and its implementation. The insights garnered from this work will inform similar studies planned for ITER[13].
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466 (PPPL), DE-FC02-04ER54698 (DIII-D), and DE-AC05-00OR22725 (ORNL). The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.

### v2

@@ 6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [33]) are shown in figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s (see poloidal wall index in figure 4a). The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-1s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-1s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10%.
Figure 6: (a) Boron surface concentrations after 400 iterations for moderate ($10^{20}$ B/s) and high ($10^{21}$ B/s) boron source strengths at the toroidal DiMES location. The horizontal broken line indicates the experimental boron concentration on DiMES. (b) XPS depth profile from a similar experiment shows 40% boron (40% carbon, with 20% oxygen and other impurities not shown) [2].
@@
Adjusting the boron injection rate modifies the surface composition of divertor targets, as shown in Figure 6. With a boron influx rate of $10^{20}$ B/s, the B surface concentrations on the inner and outer targets reach approximately 35%, with boron area densities ranging from $0.3-1.4\times 10^{16}$ cm-2. Increasing the influx rate tenfold to $10^{21}$ B/s significantly enriches the boron on the target surfaces, potentially forming nearly pure boron layers, with area densities between $1.8$ and $3.6\times 10^{16}$ cm-2. This saturation phenomenon contrasts with lower influx rates, where the surface composition sustains a mixed B-C coating, approximating a B concentration closer to experimental findings, with B and C concentrations each being 40% as shown in figure 6b. This effect can be attributed to the differential sputtering rates of B and C, which minimize the overall sputtering yields. This effect and associated boron-carbon compositions have been observed in previous DIII-D IPD experiments and traditional boronization procedures in tokamaks like TEXTOR and ASDEX-Upgrade[2, 34, 35]. For IPD experiments, however, achieving a balanced B:C concentration typically required B powder rates of 10-20 mg/s or $10^{21}$ B/s. Discussions in the literature, such as in [2], highlight a marginal dependence of surface composition on boron influx rates, underscoring the challenges of attaining pure boron layers in environments with significant carbon presence. Differences between modeling and experiment may be partially explained by the simulation domain focusing on the divertor region, with the main chamber wall not yet included. Carbon wall elements beyond the divertor might contribute to the total C influxes due to additional plasma-wall interactions (e.g., main wall CX-erosion) not considered in the present study.
@@ 7 Summary and conclusions
A new comprehensive modeling approach for real-time wall conditioning and in-situ boron powder coating is presented, utilizing three advanced models: EMC3-EIRENE, DIS, and WallDYN3D [18, 21, 25, 26, 27]. These models collectively address 3D plasma transport, powder particle migration and ablation, and mixed-material surface evolution. Ablation modeling with EMC3-EIRENE/DIS shows that boron particles of up to 150 $\mu$m evaporate at the separatrix, where they reach evaporation temperature. The transport analysis shows that initial B fluxes are higher on the inboard divertor target due to the directional frictional forces from the main ion flow in the evaporation region. This effect closely aligns with findings from EMC3-EIRENE simulations using a single atomic B source in the same plasma scenario previously [17], and is consistent with beryllium transport modeling for JET and ITER using ERO2.0 [36]. However, first-flight deposition based on EMC3-EIRENE/DIS alone neglects impurity recycling, resulting in significant asymmetries in the divertor B fluxes concentrated in a narrow, far SOL region. The powder particle mass was not found to affect the distribution of B fluxes between the inboard and outboard divertor targets for particle sizes 5-250 $\mu$m in the present plasma scenario.
@@
This integrated modeling approach provides a new framework for studying real-time wall conditioning and in-situ coating in advanced tokamak scenarios. The results suggest that poloidal and toroidal coverage of divertor plasma-facing components with boron can be achieved through single-point B powder injection within the limitations of the model and its implementation. The insights garnered from this work will inform similar studies planned for ITER[16].
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466 (PPPL), DE-FC02-04ER54698 (DIII-D), and DE-AC05-00OR22725 (ORNL). The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The data that support the findings of this study will be openly available in Princeton Data Commons at https://doi.org/10.34770/w32p-2w69.